OpenStudy (anonymous):

given b=-1 and h=-1, what is the equation of the graph if the parent function is y=sqrt x?

OpenStudy (anonymous):

A. y= sqrt x-1 B. y= sqrt -x+1 C. y= -sqrt -x-1 D. y=sqrt-x-1

OpenStudy (anonymous):

do u know answer:)?

OpenStudy (blurbendy):

i think it's D